OpenStudy (anonymous):

Really need help ASAP, would be SOOOO appreciated. Please show me how to do this, im so confused... Find the domain of the function. g(x)=(x+7)/(x^3+9x)

OpenStudy (anonymous):

find out the x values that make the denominator zero.

OpenStudy (anonymous):

you'll prolly have to factor your expression...

OpenStudy (anonymous):

so would -7, -3, 3 ...be correct? Sorry, im just terrible with math. :(

OpenStudy (anonymous):

ok... let's factor that denominator... x^3 + 9x = x(x^2 + 9) what makes that zero? let see: x(x^2+9) = 0 is it clear what x needs to be in order for x(x^2+9)=0 ?

OpenStudy (anonymous):

what can you tell me about this equation: a*b = 0

OpenStudy (anonymous):

one has to be 0

OpenStudy (anonymous):

good... that's excactly how i wrote the equation: x * (x^2+9) = 0 a * b = 0

OpenStudy (anonymous):

so either x=0 or x^2+9=0, agreed?

OpenStudy (anonymous):

yeah/... i get it now. Thanks sooooooo much. Math is my worst subject. your the best!! =D

OpenStudy (anonymous):

yw... the domain btw... is all real numbers except x=0...:)